OpenStudy (anonymous):

What is the inverse function of f(x) = 4 - x/6

OpenStudy (anonymous):

invert x and y (f(x)) we have y = 4 - x/6 invert to x = 4 - y/6 and solve for y (put it in the for form of f(x))